MPI08 Proceedings

Name1^{1}^{†}, Name2^{2}, Name3^{3} 1Institute Name1, 2Institute Name2, 3Institute Name3

Abstract

Here should be a short summary of your brilliant ideas which you will explain in more detail on the following pages

1 Your contribution

1.1 Your ingenious ideas Here comes the text, with an example citatation [1] of a famous book. And all of this is typestted with the style of cernrep supplemented with the citation option mcite, where an exapmple is shown here using the command for multiple citations [2].

1.2 How to L^{A}TEX it The references are optionally included via BiBTeX. To produce the L^{A}TEX document, do the following:

latex mpi08.tex
bibtex mpi08
latex mpi08
latex mpi08
dvips mpi08

The BiBTeX file mpi08.bib can be edited

Here comes the text, with an example citatation [1] of a famous book. And all of this is typestted with the style of cernrep supplemented with the citation option mcite, where an exapmple is shown here using the command for multiple citations [2].

References

[1] R. Feynman, Photon–Hadron Interactions. Benjamin, New York, 1972.

[2] UA5 Collaboration, G. J. Alner et al., Z. Phys. C33, 1 (1986); CDF Collaboration, F. Abe et al., Phys. Rev. D41, 2330 (1990); B. Andersson, G. Gustafson, and J. Samuelsson, Nucl. Phys. B467, 443 (1996).

†speaker